# Universal Euler-Cartan Circuits for Quantum Field Theories

### v1

@@ Conclusion.
The presented results demonstrate the power of geometric control theory [47] for the analysis of QFTs. While geometric quantum control is known to be the central concept in the search for an optimal circuit for quantum computation [48, 49, 50], previous proposal are centered around minimization of the distance in the Finsler space of unitary operators. In contrast, the current work presents an alternative framework based on the Fubini-Study metric and a universal circuit ansatz. The proposed framework is not only numerically-tractable, but also readily-implementable on current quantum machines, opening the door to hybrid quantum-classical analysis of a large class of QFTs.

### v2

@@ Conclusion.
The presented results demonstrate the power of geometric control theory [49] for the analysis of QFTs. While geometric quantum control is known to be the central concept in the search for an optimal circuit for quantum computation [50, 51, 52], previous proposal are centered around minimization of the distance in the Finsler space of unitary operators. In contrast, the current work presents an alternative framework based on the Fubini-Study metric and a universal circuit ansatz. The proposed framework is not only numerically-tractable, but also readily-implementable on current quantum simulators, enabling efficient preparation of multiple-meson or domain-wall excitations of confining QFT models in two space-time dimensions. These can shed novel insights into string fragmentation dynamics in models emulating quantum chromodynamics as well as scattering amplitudes in non-integrable perturbations of two-dimensional conformal field theories.